The bull-throated ding-kishker bird, on the planet Urepetis, shows a variable volume in its mating call. The volume is measured in Urepetian decibels (Uds). A random sample of 100 birds show the following distribution of volumes:

Volume (Uds) Frequency
0.00 - 1.00 8
1.00 - 2.00 23
2.00 - 3.00 38
3.00 - 4.00 24
4.00 - 5.00 7

Assuming that the mating call volume is a genetic trait strictly determined by underlying genes that act additively, what is the most likely number of genes that determine the mating call volume?
Answer #1

The number of genes in quantitative or polygenic inheritance is calculated by following formula :

2n+1 = number of phenotype.

Here n= number of genes

Number of phenotype in this case is 5

That's why

2n+1 = 5

2n= 5-1

2n= 4

n= 4/2

n=2.

There are two gene which determine the mating call volume.